OpenStudy (rosales0228):

Anyone love doing math and wants to help me with this pretest I will give medals and fans!!! What is the arc length of the subtending arc for an angle of 72 degrees on a circle of radius 4?

OpenStudy (triciaal):

arc length / circumference in same ratio as angle at center/360 if you let the length of the arc = l then 72/360 = l/2*pi*r given r the radius = 4

OpenStudy (rosales0228):

im confused!!!

OpenStudy (triciaal):

still waiting for your input. what are you confused about? do you understand what a ratio is? do you know that a circle has 360° and the circumference of the circle is 2*pi* r

OpenStudy (triciaal):